OpenStudy (anonymous):

Can someone help me solve (x1/4)4/9?

OpenStudy (anonymous):

\[ (x^{1/4})^{4/9} \]?

OpenStudy (anonymous):

Yep that's it! Was just trying to put up a drawing of it, but you beat me to it haha.

OpenStudy (anonymous):

(x/4)(4/9) =4x/36 =x/9

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

\[ (x^a)^b=x^{a\times b} \]So what is \(1/4\times 4/9\)?

OpenStudy (jdoe0001):

\(\huge (x^{\frac{1}{4}})^{\frac{4}{9}}\implies x^{\frac{1}{4}\times \frac{4}{9}}\)

OpenStudy (anonymous):

x^1/9